n-6 n-5
--- = ---
3n 3n+1

Answer by checkley77(12569)

n-6 n-5
--- = ---
3n 3n+1
cross multiply.
3n(n-5)=(n-6)(3n+1)
3n^2-15n=3n^2-17n-6
3n^3-3n^2-15n+17n=-6
2n=-6
n=-6/2
n=-3 ans.
Proof:
(-3-6)/-3*3=(-3-5)/(3*-3+1)
-9/-9=-8/(-9+1)
1=-8/-8
1=1